OpenStudy (anonymous):

A quantity, its 3/5, its 1/3, and its 1/4, added together, become 81. What is the quantity?

OpenStudy (anonymous):

\[\frac{3}{5}x+\frac{1}{3}x+\frac{1}{4}x=81\] \[\frac{71}{60}x=81\]

OpenStudy (anonymous):

The middle step is: x(3/5 + 1/3 + 1/4) = 81 x(36/60 + 20/60 + 15/16) = 81

OpenStudy (anonymous):

Some how the answer is 4860/131.

OpenStudy (anonymous):

Oh, wait. Satellite made a mistake. It's actually x + 3/5x + 1/3x + 1/4x = 81 x(1+3/5+1/3+1/4) = 81 x(60/60+36/60 + 20/60 + 15/60) = 81 x(131/60) = 80

OpenStudy (anonymous):

=[ I believe it's 4800/131. Not 4860...

OpenStudy (anonymous):

60(x+3/5x+1/3x+1/4x)=81 I added 60x+36x+20x+15x=4860 131x=4860 x=4860/131

OpenStudy (anonymous):

Oh haha I see my mistake. I just did 80*60 rather than 81*60.

OpenStudy (anonymous):

Its all good! I do My Math Lab for school. So that is how I figured out the set up.